A student lifts 100 N of books 1 m to his shoulder. He walks forward 10 m to the stairs and goes up 3 m. He walks 20 m forwards to class. How much work has been done to the books?
1
Expert's answer
2020-02-03T14:00:53-0500
"W=F(h+H)"

"W=100(1+3)=400\ J"